# Cascade adaptive optics with a second stage based on a Zernike wavefront sensor for exoplanet observations

### v1

@@ 5.1 Low flux regime
The ZWFS is known to be twice more sensitive sensor than the PWFS in the unmodulated configuration (Guyon 2005) since its signal projects onto a single pupil while PWFS splits the signal into four pupil images. This sensitivity gain is even larger when the comparison is made with the PWFS in modulated configuration. This advantage proves interesting in low flux regimes such as the observation of planetary companions around faint targets. To explore this possible benefit, we run an experiment with the same source flux and the same observing conditions for the ZWFS and the PWFS control loops. The parameters mag, , and are set to 5.0, 0.7” and 10 m.s-1. For both configurations, we adjust the loop gain to enable the control schemes to run without introducing noise amplification in the controlled region. The optimal integrator gain is found at 0.15 for the ZWFS and 0.03 for the PWFS in modulated configuration.
Figure 11 represents the intensity profiles of the coronagraphic images in open and closed loop for both ZWFS and PWFS control schemes. The open loop profiles exhibit differences which are related to the different static aberrations that are applied on the DM in both cases. In the PWFS case, the flat response on the sensor is based on the by-eye optimization of the coefficients related to some first low-order Zernike modes. For the ZWFS configuration, we perform a two-step calibration as detailed in Section 2.3. The wavefront pattern on the sensor results from a DM shape which produces a smooth and homogeneous intensity in the ZWFS pupil, offering a more robust control loop with this sensor, see Section A.4. The flat intensity in the ZWFS pupil corresponds to wavefront errors that are non null, leading to some residual static low-order aberrations and speckles on the science camera. In open loop, the PWFS control loop therefore offers a coronagraphic image with slightly deeper contrast than the ZWFS control loop.
@@
As this experiment presents its own limits, it is obviously risky to draw hasty conclusions and the purpose is different here. In this very preliminary test, the low flux configuration was set assuming the absence of impact on the correction of the first stage. In real life, the first-stage XAO system will see its performance limited in low flux regime and will most likely provide images with a moderate wavefront correction to the second stage AO. This situation will lead to aberration amplitudes that will certainly impact the ability of the second stage to correct for the AO residuals. In this scenario, the ZWFS control loop will be altered by the limited capture range of its sensor. Further studies therefore needs to be performed to compensate for these large residuals with the ZWFS control loop. Strategies such as the use of non-linear reconstructors (e.g. Haffert 2024) are promising options to run efficiently the ZWFS but this is out of the scope of the paper and will be addressed in future experiments.
Still, the control loop with ZWFS constitutes a complementary solution to the adopted second-stage baseline with the PWFS and further investigations can reveal the benefits of working with a control loop based on a sensor with one of the highest sensitivities.
@@ 5.2 Control loop speed
While our experiment is somehow unrealistic since the turbulence runs at 350Hz and the loop runs at 5000 Hz, the turbulence does not evolve in-between loop corrections. However, the purpose here is to show that the ZWFS control loop can possibly run at extremely high speeds. While the modulated PWFS is limited by the modulation mirror to 3 kHz, we show here that the ZWFS lifts this limitation and could be run on our off-the-shelf WFS and RTC hardware with 5 kHz. Another feature of the ZWFS is that it only observes one pupil image, so the detector area is 4 times smaller than for the PWFS and possible frame-rates are hence up to 4 times faster for a given camera depending on the readout concept.
@@ 6 Discussion and conclusions
We have presented a first proof of concept of the ZWFS-based second-stage AO loop, an alternative solution of the PWFS analog, to support a single-stage XAO system for coronagraphic observations of circumstellar environments. Our tests on GHOST validate the ability of our approach to reduce the AO residuals left by the first stage XAO loop, leading to contrast gains with a factor of at least 2 in the coronagraphic image of an observed point source. While encouraging, this gain of 2 is limited by the presence of quasi-static speckles and coronagraphic residuals in our experiment. The AO residuals are reduced by much larger factors which would pay off in terms of contrast gains with a more efficient coronagraph and NCPA calibration scheme.
@@
Our experiment shows that the ZWFS control loop provides similar contrast gains as the PWFS second-stage AO loop for the same observing conditions. As the ZWFS is twice more sensitive than the PWFS, the corresponding control loop scheme features a more robust contrast gain in low-flux regime, proving interesting for the observation of planetary companions around faint nearby stars. The PWFS can operate in modulated mode to improve the wavefront error correction of its control loop. This configuration presents a maximum speed (e.g. 3 kHz for the tip-tilt mirror modulation in the envisioned VLT/SPHERE+ upgrade Boccaletti et al. 2022) which limits the AO correction frequency of the PWFS-based control loop in modulated mode. The ZWFS control loop is free of modulation aspects, enabling wavefront control at speeds of up to 5 kHz or higher depending on the WFS camera and RTC hardware, provided that the natural guide star is bright enough for wavefront sensing. Such correction rates pave the way for wavefront correction in very high wind speed conditions and therefore, a possible extension of the useful observing time and science return for the observation of exoplanetary systems.
@@
In our experiment, the ultimate contrast is mainly limited by the diffraction residuals from the classical Lyot coronagraph and by the NCPA existing between the control loop and the science path in the testbed. The first limitation can be removed by upgrading the starlight suppression system with a more elaborated coronagraph design to provide deeper contrast (e.g. reviews by Guyon et al. 2006; Ruane et al. 2018). The second limitation can be addressed by estimating the aberrations in science path, using e.g. some non invasive strategies (e.g. phase diversity, differential optical transfer function) to calibrate the static errors, and then modifying the reference slopes in the calibration matrix. A preliminary estimate of the aberrations was already performed with the Fast & Furious phase diversity algorithm in this paper. Additional real-time control strategies might be required to compensate for the quasi-static aberrations in GHOST.
@@
The validation of our ZWFS-based approach is performed in monochromatic light. In real-life facilities, wide spectral band filters are often used to maximize the amount of photons on the WFS and improve the wavefront correction. Theoretical studies have already showed the efficiency of the ZWFS for measuring aberrations in broadband light (N’Diaye et al. 2013; Haffert 2024). In the forthcoming tests, we plan to demonstrate the capability of ZWFS control loop to work in broadband light by installing a white light source and filters on GHOST.
@@
In adaptive optics systems, photon sharing between the wavefront sensing path and the science path often requires an adequate balance between both arms to ensure an optimal wavefront error correction without jeopardizing the observation of the faintest substellar mass companions. While the ZWFS is well-known for its high sensitivity, recent approaches have emerged to increase the sensor sensitivity even further by adjusting the mask size (Chambouleyron et al. 2021) or by combining the sensor with an appropriate entrance pupil apodization (Haffert et al. 2023). These compelling approaches can push the wavefront correction further for the ZWFS control loop, making it an even more attractive solution for a two-stage AO loop.
@@
The ZWFS control loop relies on the linear capture range of the sensor to correct for the wavefront errors left from the first XAO stage. Since the ZWFS is renowned for its limited capture range, new flavors have recently emerged for this sensor to extend its capture range of the sensor further, such as the vector ZWFS (Doelman et al. 2019), the phase-shifting ZWFS (Wallace et al. 2011), the combination of the ZWFS with phase diversity algorithms (Haffert 2024), or the use of the ZWFS in multi-wavelength approach (Vigan et al. 2011; Haffert 2024). Non-linear reconstructors based on neural network also constitute another promising avenue to extend the capture range of our approach, as it was recently proposed with the unmodulated PWFS (Landman et al. 2024). These schemes offer promising prospects to extend the sensor capture range further and broaden the use of the ZWFS wavefront control.
@@
In our study, the ZWFS control loop works in standalone mode, meaning the absence of interaction between the loops of the two AO stages. Command laws integrating the information of both loops represent a rising path to derive new approaches for the two-stage AO loop and maximize the overall correction of the wavefront errors due to the atmospheric turbulence. In this context, the temporal controller constitutes another key aspect to investigate further for the ZWFS control loop. Our current study relies on the use of an integrator in standalone mode. In future works, we will investigate more advanced command laws based on predictive control algorithms (Poyneer et al. 2007; Males & Guyon 2018) or reinforcement learning methods (Nousiainen et al. 2024) to further enhance the wavefront correction regime, the speed and accuracy of the ZWFS control loop in standalone and in an integrated version with the first AO stage.
@@
Finally, on-sky demonstration is a crucial step to increase the readiness level of our approach. Some on-sky studies with the Provence Adaptive-optics PYramid RUn System (PAPYRUS) testbed at Observatoire de Haute Provence in France are currently on going to demonstrate the feasibility of the two-stage AO with a vector ZWFS in the second loop (Cisse et al. 2023). The expected VLT/SPHERE+ project serves as a technology demonstrator to prepare for XAO facilities with ELTs for exoplanet imaging and spectroscopy (Boccaletti et al. 2022). The PWFS is selected in the baseline for the second-stage AO loop in the SPHERE upgrade. We are currently investigating the implementation of a ZWFS as an additional option to the PWFS to enlarge the wavefront correction options, explore suitable operating modes that could be left uncovered by the PWFS option, and maximize the science return for exoplanet imagers with ELTs. The exploration and in-lab validation of multiple and complementary second-stage AO strategies will be an asset for the most challenging observations of exoplanets from the ground, such as Earth-like planets around M dwarfs with the ELTs (Kasper et al. 2021; Fitzgerald et al. 2022; Kautz et al. 2023).
@@ Acknowledgements.
This work was supported by the Action Spécifique Haute Résolution Angulaire (ASHRA) of CNRS/INSU co-funded by CNES. M. N’Diaye acknowledges support from Observatoire de la Côte d’Azur and Laboratoire Lagrange through the 2022 BQR OCA and 2023 BQR Lagrange programs for the manufacturing of Zernike phase masks and the missions to ESO Garching. AV acknowledges funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme, grant agreement No. 757561 (HiRISE).References
@@ A.1 Number of corrected KL modes
We analyze the effect of the number of corrected KL modes on the ZWFS control loop behavior. Figure 12 shows the intensity profiles of the coronagraphic images in open and closed loop for different . As expected, the ZWFS-based controlled region increases with . Beyond 350 KL modes, no further gain is observed, showing limitations in terms of wavefront correction. The DM contains 24 actuators across the pupil diameter, allowing us to control wavefront errors with spatial frequencies up to 12 cycles/pupil (c/pup) or equivalently up 122, i.e. 452 KL modes.
@@ A.2 Control loop gain
We investigate the impact of the control loop gain on the ZWFS-based second-stage AO performance. Figure 13 displays the intensity profiles of the coronagraphic images in open and closed loops for different integrator gains. There is clearly no contrast difference between all the considered loop gains. As the ZWFS control loop runs twice faster than the first stage XAO loop, temporal error does not represent a limitation in our experiment and hence, the loop gain has no incidence in the wavefront correction and image contrast. These results are achieved with median observing conditions. In less favorable atmospheric turbulence, the loop gain remains a free parameter to adjust to allow for noisy WFS measurements to be averaged and reduce photon noise.
@@ A.3 Field stop
The field stop enables to reduce the aliasing effects in the ZWFS measurements after the first XAO stage by adjusting the pinhole diameter. We study the influence of the field stop size on the behavior of our ZWFS control loop. Figure 14 represents the intensity profiles of the coronagraphic images in open and closed loops for two field stop sizes. As expected, the reduction of the field stop size improves the control loop performance with a contrast gain going from 1.5 to 2 in median observing conditions. This result confirms the benefit of using an adjustable field stop in the second stage AO loop to reduce aliasing effects. The size of the field stop can be tuned depending on the atmospheric turbulence conditions. In the core of the paper, we keep on working with the smallest field stop with a diameter of 35 .
@@ A.4 Control loop calibration
Figure 15 displays the intensity profiles of the coronagraphic images for both calibrations. The ZWFS control loop works in both cases with similar contrast gains, showing its effectiveness for both calibration schemes. During our experiments, our control loop proves more stable when working around the ZWFS flat map. The ZWFS is known for having an asymmetric capture range with a small linear range (e.g. N’Diaye et al. 2013). The control loop with the ZWFS benefits from setting the reference slopes to a sensor response flat map since it provides a uniform response at which the ZWFS can properly measure wavefront errors and maintain a suitable working control loop. However, as the sensor flat response is different from the response to zero aberration as recalled in Eq. (2), it results in the introduction of a static aberration in our control loop.
Shifting the reference slopes from the ZWFS to the DM flat map enables to work with the system to achieve null phase errors. This comes at a cost since this calibration relies on the asymmetry in the capture range of the sensor for each subaperture. It results in higher risks to get out from the linear regime of the sensor, leading to situations in which the control loop can break down and open. The ZWFS control loop efficiently works in different calibration situations and its performance is the most stable for the calibration around the ZWFS flat map.

### v2

@@ 5.1 Low flux regime
The ZWFS is known to be twice more sensitive sensor than the PWFS in the unmodulated configuration (Guyon 2005) since its signal projects onto a single pupil, while PWFS splits the signal into four pupil images. This sensitivity gain is even larger when the comparison is made with the PWFS in modulated configuration. This advantage proves interesting in low flux regimes such as the observation of planetary companions around faint targets. To explore this possible benefit, we run an experiment with the same source flux and the same observing conditions for the ZWFS and the PWFS control loops. The parameters mag, , and were set to 5.0, 0.7”, and 10 m.s-1. For both configurations, we adjusted the loop gain to enable the control schemes to run without introducing noise amplification in the controlled region. The optimal integrator gain is found at 0.15 for the ZWFS and 0.03 for the PWFS in modulated configuration.
Figure 11 represents the intensity profiles of the coronagraphic images in open and closed loop for both ZWFS and PWFS control schemes. The open loop profiles exhibit differences which are related to the different static aberrations that are applied on the DM in both cases. In the PWFS case, the flat response on the sensor is based on the by-eye optimization of the coefficients related to some first low-order Zernike modes. For the ZWFS configuration, we perform a two-step calibration as detailed in Sect. 2.3. The wavefront pattern on the sensor results from a DM shape which produces a smooth and homogeneous intensity in the ZWFS pupil, offering a more robust control loop with this sensor, see Sect. A.4. The flat intensity in the ZWFS pupil corresponds to wavefront errors that are non null, leading to some residual static low-order aberrations and speckles on the science camera. In an open loop, the PWFS control loop therefore offers a coronagraphic image with slightly deeper contrast than the ZWFS control loop.
@@
As this experiment presents its own limits, it is obviously risky to draw hasty conclusions. Thus, in this very preliminary test, the low flux configuration was set assuming the absence of impact on the correction of the first stage. In real life, the first-stage XAO system will see its performance limited in low flux regime and will most likely provide images with a moderate wavefront correction to the second stage AO. This situation will lead to aberration amplitudes that will certainly impact the ability of the second stage to correct for the AO residuals. In this scenario, the ZWFS control loop will be altered by the limited capture range of its sensor. Further studies therefore needs to be performed to compensate for these large residuals with the ZWFS control loop. Strategies such as the use of non-linear reconstructors (e.g., Haffert 2024) are promising options to run efficiently the ZWFS but this is out of the scope of the paper and will be addressed in future experiments.
Still, the control loop with ZWFS constitutes a complementary solution to the adopted second-stage baseline with the PWFS. Further investigations can reveal the benefits of working with a control loop based on a sensor with one of the highest sensitivities.
@@ 5.2 Control loop speed
While our experiment is somehow unrealistic since the turbulence runs at 350Hz and the loop runs at 5000 Hz, the turbulence does not evolve in-between loop corrections. However, the purpose here is to show that the ZWFS control loop can possibly run at extremely high speeds. While the modulated PWFS is limited by the modulation mirror to 3 kHz, we show here that the ZWFS lifts this limitation and could be run on our off-the-shelf WFS and RTC hardware with 5 kHz. Another feature of the ZWFS is that it only observes one pupil image, so the detector area is four times smaller than for the PWFS. Then, possible frame-rates are up to four times faster for a given camera depending on the readout concept.
@@ 6 Discussion and conclusions
We present a first proof of concept for the ZWFS-based second-stage AO loop, an alternative solution of the PWFS analog, to support a single-stage XAO system for coronagraphic observations of circumstellar environments. Our tests on GHOST validate the ability of our approach to reduce the AO residuals left by the first stage XAO loop, leading to contrast gains with a factor of at least 2 in the coronagraphic image of an observed point source. While encouraging, this gain of 2 is limited by the presence of quasi-static speckles and coronagraphic residuals in our experiment. The AO residuals are reduced by much larger factors, which would pay off in terms of contrast gains with a more efficient coronagraph and NCPA calibration scheme.
@@
Our experiment shows that the ZWFS control loop provides similar contrast gains as the PWFS second-stage AO loop for the same observing conditions. As the ZWFS is twice more sensitive than the PWFS, the corresponding control loop scheme features a more robust contrast gain in low-flux regime, proving interesting for the observation of planetary companions around faint nearby stars. The PWFS can operate in modulated mode to improve the wavefront error correction of its control loop. This configuration presents a maximum speed (e.g., 3 kHz for the tip-tilt mirror modulation in the envisioned VLT/SPHERE+ upgrade Boccaletti et al. 2022), which limits the AO correction frequency of the PWFS-based control loop in modulated mode. The ZWFS control loop is free of modulation aspects, enabling wavefront control at speeds of up to 5 kHz or higher depending on the WFS camera and RTC hardware, provided that the natural guide star is bright enough for wavefront sensing. Such correction rates pave the way for wavefront correction in very high wind speed conditions and, therefore, a possible extension of the useful observing time and science return for the observation of exoplanetary systems.
@@
In our experiment, the ultimate contrast is mainly limited by the diffraction residuals from the classical Lyot coronagraph and by the NCPA existing between the control loop and the science path in the testbed. The first limitation can be removed by upgrading the starlight suppression system with a more elaborated coronagraph design to provide deeper contrast (see, e.g., reviews by Guyon et al. 2006; Ruane et al. 2018). The second limitation can be addressed by estimating the aberrations in science path, for instance, by using some non-invasive strategies (e.g., phase diversity, differential optical transfer function) to calibrate the static errors, and then modifying the reference slopes in the calibration matrix. A preliminary estimate of the aberrations was already performed with the Fast & Furious phase diversity algorithm in this paper. Additional real-time control strategies might be required to compensate for the quasi-static aberrations in GHOST.
@@
The validation of our ZWFS-based approach is performed in monochromatic light. In real-life facilities, wide spectral band filters are often used to maximize the amount of photons on the WFS and improve the wavefront correction. Theoretical studies have already demonstrated the efficiency of the ZWFS for measuring aberrations in broadband light (N’Diaye et al. 2013; Haffert 2024). In the forthcoming tests, we plan to demonstrate the capability of ZWFS control loop to work in broadband light by installing a white light source and filters on GHOST.
@@
In AO systems, photon sharing between the wavefront sensing path and the science path often requires an adequate balance between both arms to ensure an optimal wavefront error correction without jeopardizing the observation of the faintest substellar mass companions. While the ZWFS is well-known for its high sensitivity, recent approaches have emerged to increase the sensor sensitivity even further by adjusting the mask size (Chambouleyron et al. 2021) or by combining the sensor with an appropriate entrance pupil apodization (Haffert et al. 2023). These compelling approaches can push the wavefront correction further for the ZWFS control loop, making it an even more attractive solution for a two-stage AO loop.
@@
The ZWFS control loop relies on the linear capture range of the sensor to correct for the wavefront errors left from the first XAO stage. Since the ZWFS is renowned for its limited capture range, new flavors have recently emerged for this sensor to extend its capture range of the sensor further. These include the vector ZWFS (Doelman et al. 2019), phase-shifting ZWFS (Wallace et al. 2011), combination of the ZWFS with phase diversity algorithms (Haffert 2024), or the use of the ZWFS in multi-wavelength approach (Vigan et al. 2011; Haffert 2024). Non-linear reconstructors based on neural network also constitute another promising avenue to extend the capture range of our approach, as it was recently proposed with the unmodulated PWFS (Landman et al. 2024). These schemes offer promising prospects to extend the sensor capture range further and broaden the use of the ZWFS wavefront control.
@@
In our study, the ZWFS control loop works in standalone mode, meaning the absence of interaction between the loops of the two AO stages. Command laws integrating the information of both loops represent a rising path to derive new approaches for the two-stage AO loop and maximize the overall correction of the wavefront errors due to the atmospheric turbulence. In this context, the temporal controller constitutes another key aspect to investigate further for the ZWFS control loop. Our current study relies on the use of an integrator in standalone mode. In future works, we will investigate more advanced command laws based on predictive control algorithms (Poyneer et al. 2007; Males & Guyon 2018) or reinforcement learning methods (Nousiainen et al. 2024) to further enhance the wavefront correction regime, along with the speed and accuracy of the ZWFS control loop in a standalone and in an integrated version with the first AO stage.
@@
Finally, on-sky demonstration is a crucial step to increase the readiness level of our approach. Some on-sky studies with the Provence Adaptive-optics PYramid RUn System (PAPYRUS) testbed at Observatoire de Haute Provence in France are currently on going to demonstrate the feasibility of the two-stage AO with a vector ZWFS in the second loop (Cisse et al. 2023). The expected VLT/SPHERE+ project serves as a technology demonstrator to prepare for XAO facilities with ELTs for exoplanet imaging and spectroscopy (Boccaletti et al. 2022). The PWFS is selected in the baseline for the second-stage AO loop in the SPHERE upgrade. We are currently investigating the implementation of a ZWFS as an additional option to the PWFS to enlarge the wavefront correction options, while exploring suitable operating modes that could be left uncovered by the PWFS option and maximizing the science return for exoplanet imagers with ELTs. The exploration and in-lab validation of multiple and complementary second-stage AO strategies will be an asset for the most challenging observations of exoplanets from the ground, such as Earth-like planets around M dwarfs with ELTs (Kasper et al. 2021; Fitzgerald et al. 2022; Kautz et al. 2023).
@@ Acknowledgements.
This work was supported by the Action Spécifique Haute Résolution Angulaire (ASHRA) of CNRS/INSU co-funded by CNES. MN acknowledges support from Observatoire de la Côte d’Azur and Laboratoire Lagrange through the 2022 BQR OCA and 2023 BQR Lagrange programs for the manufacturing of Zernike phase masks and the missions to ESO Garching. AV acknowledges funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme, grant agreement No. 757561 (HiRISE).References
@@ A.1 Number of corrected KL modes
We analyze the effect of the number of corrected KL modes on the ZWFS control loop behavior. Figure 12 shows the intensity profiles of the coronagraphic images in open and closed loop for different . As expected, the ZWFS-based controlled region increases with . Beyond 350 KL modes, no further gain is observed, showing limitations in terms of wavefront correction. The DM contains 24 actuators across the pupil diameter, allowing us to control wavefront errors with spatial frequencies up to 12 cycles/pupil (c/pup) or equivalently up to 122, namely, 452 KL modes.
@@ A.2 Control loop gain
We investigated the impact of the control loop gain on the ZWFS-based second-stage AO performance. Figure 13 displays the intensity profiles of the coronagraphic images in open and closed loops for different integrator gains. There is clearly no contrast difference between all the considered loop gains. As the ZWFS control loop runs twice faster than the first stage XAO loop, temporal error does not represent a limitation in our experiment and hence, the loop gain has no incidence in the wavefront correction and image contrast. These results are achieved with median observing conditions. In less favorable atmospheric turbulence, the loop gain remains a free parameter to adjust to allow for noisy WFS measurements to be averaged and reduce photon noise.
@@ A.3 Field stop
The field stop allows us to reduce the aliasing effects in the ZWFS measurements after the first XAO stage by adjusting the pinhole diameter. We study the influence of the field stop size on the behavior of our ZWFS control loop. Figure 14 represents the intensity profiles of the coronagraphic images in open and closed loops for two field stop sizes. As expected, the reduction of the field stop size improves the control loop performance with a contrast gain going from 1.5 to 2 in median observing conditions. This result confirms the benefit of using an adjustable field stop in the second stage AO loop to reduce aliasing effects. The size of the field stop can be tuned depending on the atmospheric turbulence conditions. In the core of the paper, we keep on working with the smallest field stop with a diameter of 35 .
@@ A.4 Control loop calibration
Figure 15 displays the intensity profiles of the coronagraphic images for both calibrations. The ZWFS control loop works in both cases with similar contrast gains, showing its effectiveness for both calibration schemes. During our experiments, our control loop proves more stable when working around the ZWFS flat map. The ZWFS is known for having an asymmetric capture range with a small linear range (e.g., N’Diaye et al. 2013). The control loop with the ZWFS benefits from setting the reference slopes to a sensor response flat map since it provides a uniform response at which the ZWFS can properly measure wavefront errors and maintain a suitable working control loop. However, as the sensor flat response is different from the response to zero aberration as recalled in Eq. (2), it results in the introduction of a static aberration in our control loop.
Shifting the reference slopes from the ZWFS to the DM flat map enables us to work with the system to achieve null phase errors. This comes at a cost since this calibration relies on the asymmetry in the capture range of the sensor for each subaperture. It results in higher risks when moving away from the linear regime of the sensor, leading to situations in which the control loop can break down and open. The ZWFS control loop efficiently works in different calibration situations and its performance is the most stable for the calibration around the ZWFS flat map.